# Easy3E: Feed-Forward 3D Asset Editing via Rectified Voxel Flow

### v1

@@ Abstract
Existing 3D editing methods rely on computationally intensive scene-by-scene iterative optimization and suffer from multi-view inconsistency. We propose an effective and fully feedforward 3D editing framework based on the TRELLIS generative backbone, capable of modifying 3D models from a single editing view. Our framework addresses two key issues: adapting training-free 2D editing to structured 3D representations, and overcoming the bottleneck of appearance fidelity in compressed 3D features. To ensure geometric consistency, we introduce Voxel FlowEdit, an edit-driven flow in the sparse voxel latent space that achieves globally consistent 3D deformation in a single pass. To restore high-fidelity details, we develop a normal-guided single to multi-view generation module as an external appearance prior, successfully recovering high-frequency textures. Experiments demonstrate that our method enables fast, globally consistent, and high-fidelity 3D model editing.
@@ 1 Introduction
To address these challenges, we propose a fully feed-forward 3D editing framework built on the TRELLIS generative backbone [45]. Given a single edited view and a user-defined editable region, our method performs both geometric and appearance editing directly in TRELLIS’s sparse voxel latent space. We introduce Voxel FlowEdit, a latent-space editing mechanism that translates source voxel to target voxel via an adapted velocity field, enabling globally coherent geometric deformation in a single pass. Following this coarse transformation, we apply a structured latent repainting stage to locally refine geometry and appearance while anchoring unedited regions, ensuring consistency and detail preservation.
@@
  • We introduce Voxel FlowEdit, a voxel-flow editing mechanism. It constructs the translation of source/target 3D assets within the sparse voxel latent space by utilizing a specially adapted velocity field, achieving globally coherent 3D geometric deformation.
@@ 2 Related Work
Recent progress in 3D generative modeling has evolved from lifting 2D observations into 3D structures [34, 22] to learning fully 3D-native representations that jointly model geometry and appearance [14]. Early image-to-3D frameworks employed NeRF or mesh decoders to reconstruct textured assets from single or few views [29, 32, 53, 33], while subsequent diffusion-based pipelines further leveraged pretrained 2D priors for text-to-3D synthesis via score distillation [34, 43, 22, 36]. To improve view consistency and scalability, several works proposed generating multi-view images as intermediate supervision before reconstructing 3D assets [26, 48, 39, 25], achieving better alignment but still constrained by 2D lifting. More recently, large-scale 3D-native frameworks have emerged, learning structured latent spaces or volumetric primitives directly from massive 3D corpora [14, 40, 47, 54, 45, 21]. These models enable feed-forward generation of meshes, radiance fields, or 3D Gaussian representations conditioned on images or text, substantially advancing fidelity and controllability. Our study builds upon this trajectory, focusing on efficient and consistent 3D editing within a structured latent space.
@@
Early diffusion-based image editing typically reconstructs a given image by inverting it into the latent space of a pretrained model and then applying localized manipulations via attention control, prompt/mask guidance, or latent blending to realize semantic and structural changes [12, 16, 28, 30, 4, 46, 41, 49]. In parallel, training-based approaches adapt the generator or lightweight adapters to the target domain (e.g., identity or structure), improving edit fidelity and controllability through finetuning or conditioning modules such as DreamBooth[37, 19], LoRA[15], and ControlNet [51, 10]. In contrast to diffusion-style U-Net editors, inversion-free flow-matching formulations directly construct continuous source–target transformations in the learned velocity field, avoiding iterative inversion and aligning naturally with DiT-style architectures [24, 18, 31]. Given that paired 3D editing data are largely unavailable and that the underlying generative backbone adopts a flow-matching formulation [24, 27], this compatibility makes the flow-based editing paradigm particularly well suited to our setting.
@@
Recent 3D editing methods can be broadly characterized by how 2D guidance is coupled to the 3D representation. A first line iteratively optimizes a 3D representation by supervising rendered views with images edited by pretrained 2D models (e.g., text or mask conditioned), typically via score-distillation losses or gradient guidance [11, 55, 38]. Building on this setup, subsequent work improves robustness and efficiency by synchronizing multi-view constraints or imposing geometry-aware priors during optimization [2, 44, 3]. A parallel direction leverages multi-view or video diffusion to produce edited view sets with stronger viewpoint coverage and spatial–temporal regularization before lifting back to 3D, enabling multi-view propagation of edits [6]. More recently, 3D-native generative backbones have begun to explore feed-forward editing. These approaches typically rely on local mechanisms such as masked reconstruction [8] or multi-view inpainting [1] to localize changes. While these methods improve efficiency, their reliance on local masking or inpainting often limits them to textural changes or simple additions, struggling with edits that require globally coherent geometric deformation. This highlights a critical need for a new paradigm that can propagate complex edits throughout the 3D latent space in a single pass.
@@ 3.2 Sparse Voxel Editing
$\mathcal{E}_{\text{sil}}(\mathbf{x})$ $=\mathrm{BCE}\!\big(S(\mathbf{x}),\,M_{\text{sil}}\big),$ (3)
@@ 3.3 SLAT Repainting
$\quad+(1-\mathcal{M}_{z})\odot\Big[(1-t_{k})\,\mathbf{z}^{\text{src}}+t_{k}\,\boldsymbol{\epsilon}_{k}\Big],$
@@ 4.1 Implementation Details
We construct our evaluation set using 100 3D assets collected from Sketchfab Website, the NPHM [9] dataset for real human heads, the THuman2.1 [50] dataset for clothed human bodies, and the Objaverse [7] dataset for objects. These datasets span a wide range of object categories and editing scenarios, allowing us to assess both stylized and photorealistic edits.
@@ 4.2 Comparisons
Figure 4: Qualitative comparison. Our method achieves clean geometry and consistent appearance across multiple views, faithfully realizing the target edits while preserving unedited regions. Competing methods either retain the original geometry (MVEdit) or exhibit strong structural distortion and inconsistency (Vox-E, Instant3dit).
We compare our method against several representative works. (1) TRELLIS [45], which serves as our generative backbone. We feed the edited target image $I_{\text{tgt}}$ directly into its 3D generation pipeline and evaluate its ability to reproduce the edit while preserving consistency with the original 3D model. (2) MVEdit [6], a multi-view diffusion framework that uses a training-free 3D adapter to jointly denoise rendered views and output textured meshes. (3) Vox-E [38], a voxel-based volumetric editing method that learns a volumetric representation from oriented images and edits existing 3D objects under diffusion priors. (4) Instant3dit [1], a recent feed-forward multiview inpainting framework for fast editing of 3D content by generating consistent views and reconstructing the edited result.
Since methods (2)(4)(2)–(4) are primarily text-guided frameworks rather than direct image-driven 3D generation, we provide them with a unified text prompt that semantically describes our image-based edit to ensure a fair comparison.
Qualitative Comparison. As shown in Fig. 4, our method produces clean, view-consistent edits that accurately realize the target modifications while preserving the global structure of the source asset. Compared to TRELLIS, which reconstructs directly from the edited view, our framework better maintains identity consistency with the original asset and avoids overfitting to a single observation. Multi-view diffusion models such as MVEdit can generate plausible textures, but in our evaluation they usually introduce negligible geometric change and suffer from view-wise inconsistency. Both Vox-E and Instant3dit fail to maintain structural integrity and multi-view consistency under complex edits. Among them, Vox-E requires significantly longer inference time, while Instant3dit runs faster but still struggles to produce stable, semantically aligned geometric edits.
In contrast, our approach achieves geometrically stable and visually realistic results within a feed-forward pass. Notably, the mask-aware latent editing preserves unedited regions, while the additional refinement stage enhances local detail and lighting continuity. These results demonstrate that our framework provides a favorable balance among edit controllability, identity preservation, and visual fidelity.
Quantitative Comparison. We evaluate performance using four widely adopted quantitative metrics: CLIP-T (text–image alignment) [35], DINO-I (perceptual quality) [5], LPIPS (perceptual similarity) [52], and FID (distributional realism) [13]. As shown in Tab. 1, our method consistently achieves the best performance across all four metrics. These results confirm that our framework produces higher overall 3D quality and text alignment than representative baselines, yielding more coherent, high-fidelity 3D assets.

### v2

@@ Abstract
Existing 3D editing methods rely on computationally intensive scene-by-scene iterative optimization and suffer from multi-view inconsistency. We propose an effective and feed-forward 3D editing framework based on the TRELLIS generative backbone, capable of modifying 3D models from a single editing view. Our framework addresses two key issues: adapting training-free 2D editing to structured 3D representations, and overcoming the bottleneck of appearance fidelity in compressed 3D features. To ensure geometric consistency, we introduce Voxel FlowEdit, an edit-driven flow in the sparse voxel latent space that achieves globally consistent 3D deformation in a single pass. To restore high-fidelity details, we develop a normal-guided single to multi-view generation module as an external appearance prior, successfully recovering high-frequency textures. Experiments demonstrate that our method enables fast, globally consistent, and high-fidelity 3D model editing.
@@ 1 Introduction
To address these challenges, we propose a fully feed-forward 3D editing framework built on the TRELLIS generative backbone [48]. Given a single edited view and a user-defined editable region, our method performs both geometric and appearance editing directly in TRELLIS’s sparse voxel latent space. We introduce Voxel FlowEdit, a latent-space editing mechanism that translates source voxels to target voxels via an adapted velocity field, enabling globally coherent geometric deformation in a single pass. Following this coarse transformation, we apply a structured latent repainting stage to locally refine geometry and appearance while anchoring unedited regions, ensuring consistency and detail preservation.
@@
  • We introduce Voxel FlowEdit, a voxel-flow editing mechanism. It constructs the source-to-target translation of 3D assets within the sparse voxel latent space by utilizing a specially adapted velocity field, achieving globally coherent 3D geometric deformation.
@@ 2 Related Work
Recent progress in 3D generative modeling has evolved from lifting 2D observations [36, 24] to learning fully 3D-native representations that jointly model geometry and appearance [15]. Early image-to-3D frameworks employed NeRF or mesh decoders to reconstruct assets from few views [31, 34, 57, 35], while subsequent diffusion-based pipelines leveraged 2D priors for text-to-3D synthesis via score distillation [36, 45, 24, 38]. To improve view consistency and scalability, several works proposed generating multi-view images as intermediate supervision before reconstructing 3D assets [28, 51, 41, 27], achieving better alignment but still constrained by 2D lifting. More recently, large-scale 3D-native frameworks have emerged, learning structured latent spaces directly from massive 3D corpora [15, 42, 50, 58, 48, 23]. These models enable feed-forward generation of meshes, radiance fields, or 3D Gaussian representations conditioned on images or text, substantially advancing fidelity and controllability. Parallelly, autoregressive models like SAR3D [7] utilize multi-scale 3D VQVAE for unified generation and understanding, while OctGPT [46] leverages octree-based representations to scale up 3D shape synthesis.
@@
Early diffusion-based image editing reconstructs a given image by inverting it into the latent space of a pretrained model and then applying localized manipulations via attention control or latent blending to realize semantic and structural changes [13, 17, 30, 32, 4, 49, 43, 53]. In parallel, training-based approaches adapt the generator or lightweight adapters to the target domain, improving edit fidelity and controllability through finetuning or conditioning modules such as DreamBooth [39, 20], LoRA [16], and ControlNet [55, 11]. In contrast to diffusion-style U-Net editors, inversion-free flow-matching formulations directly construct continuous source-target transformations in the learned velocity field, avoiding iterative inversion and aligning naturally with DiT-style architectures [26, 19, 33]. Given that paired 3D editing data are largely unavailable and that the underlying generative backbone adopts a flow-matching formulation [26, 29], this compatibility makes the flow-based editing paradigm well suited to our setting.
@@
Recent 3D editing methods are characterized by how 2D guidance is coupled to the 3D representation. A first line iteratively optimizes 3D representations by supervising rendered views with 2D-edited images, typically via score-distillation losses [12, 59, 40]. Subsequent work improves robustness and efficiency by synchronizing multi-view constraints or imposing geometry-aware priors during optimization [2, 47, 3]. A parallel direction leverages multi-view or video diffusion to produce edited view sets with stronger viewpoint coverage and spatial-temporal regularization before lifting back to 3D, enabling multi-view propagation of edits [6]. More recently, 3D-native generative backbones have begun to explore feed-forward editing. These approaches typically rely on local mechanisms such as masked reconstruction [9] or multi-view inpainting [1] to localize changes. While these methods improve efficiency, their reliance on local masking or inpainting often limits them to textural changes or simple additions, struggling with edits that require globally coherent geometric deformation. Contemporaneous to our work, VoxHammer [21] and Nano3D [52] independently explore a highly similar pipeline for feed-forward editing on 3D-native generative models. Developed in parallel, our study shares the core motivation of enabling direct manipulations within the 3D latent space. These works collectively demonstrate the timeliness and significance of shifting towards fully 3D-native feed-forward editing paradigms.
@@ 3.2 Sparse Voxel Editing
$\mathcal{E}_{\text{sil}}(\mathbf{x})$ $=\operatorname{BCE}\!\big(S(\mathbf{x}),\,M_{\text{sil}}\big),$ (3)
@@ 3.3 SLAT Repainting
$\quad+(1-\mathcal{M}_{z})\odot\Big[(1-t_{k-1})\,\mathbf{z}^{\text{src}}+t_{k-1}\,\boldsymbol{\epsilon}_{k-1}\Big],$
@@ 4.1 Implementation Details
We construct our evaluation set using 100 3D assets collected from Sketchfab Website, the NPHM [10] dataset for real human heads, the THuman2.1 [54] dataset for clothed human bodies, and the Objaverse [8] dataset for objects.
@@ 4.2 Comparisons
Figure 4: Qualitative comparison. Our method achieves clean geometry and consistent appearance across multiple views, faithfully realizing the target edits while strictly preserving unedited regions. Competing methods either retain the original geometry (MVEdit) or exhibit strong structural distortion and inconsistency (Vox-E, Instant3dit). Although VoxHammer also utilizes mask inputs to guide the editing process, it struggles to maintain high fidelity outside the masked areas.
We compare our method against several representative works. (1) TRELLIS [48], which serves as our generative backbone. We feed the edited target image $I_{\text{tgt}}$ directly into its 3D generation pipeline and evaluate its ability to reproduce the edit while preserving consistency with the original 3D model. (2) MVEdit [6], a multi-view diffusion framework that uses a training-free 3D adapter to jointly denoise rendered views and output textured meshes. (3) Vox-E [40], a voxel-based volumetric editing method that learns a volumetric representation from oriented images and edits existing 3D objects under diffusion priors. (4) Instant3dit [1], a feed-forward multiview inpainting framework for fast 3D editing via view generation and reconstruction. (5) VoxHammer [21], a feed-forward 3D editing framework built upon TRELLIS that uses an image and a mask to guide the editing process. Since methods (2)-(4) are primarily text-guided frameworks rather than direct image-driven 3D generation, we provide them with a unified text prompt that semantically describes our image-based edit to ensure a fair comparison.
Qualitative Comparison. As shown in Fig. 4, our method produces clean, view-consistent edits while preserving the source asset’s global structure. Compared to TRELLIS and VoxHammer, our framework avoids overfitting and better maintains identity consistency with the original 3D model. MVEdit generates plausible textures but introduces negligible geometric changes and view-wise inconsistency. Both Vox-E and Instant3dit fail to maintain structural integrity under complex edits, with Vox-E also requiring significantly longer inference time. Overall, our approach achieves a superior balance of edit controllability, identity preservation, and visual fidelity.
Quantitative Comparison. We evaluate performance using four widely adopted quantitative metrics: CLIP-T (text-image alignment) [37], DINO-I (perceptual quality) [5], LPIPS (perceptual similarity) [56], and FID (distributional realism) [14]. As shown in Tab. 1, our method consistently achieves the best performance across all four metrics. These results confirm that our framework produces higher overall 3D quality and text alignment than representative baselines, yielding more coherent, high-fidelity 3D assets.